### v1

# Prot2Chat: Protein LLM with Early Fusion of Sequence and Structure

Abstract
Proteins play a pivotal role in living organisms, yet understanding their functions presents significant challenges, including the limited flexibility of classification-based methods, the inability to effectively leverage spatial structural information, and the lack of systematic evaluation metrics for protein Q&A systems. To address these limitations, we propose Prot2Chat, a novel framework that integrates multimodal protein representations with natural language through a unified module, enabling large language model (LLM)-driven answer generation. Our model incorporates a modified ProteinMPNN encoder, which encodes protein sequence and structural information in a unified manner, a protein-text adapter with cross-attention mechanisms, and a LLaMA3 decoder. To optimize training efficiency, we freeze the encoder and employ LoRA techniques for the decoder. We conducted experiments on two datasets, both automated metrics and expert evaluations demonstrate the superior performance of our model. Furthermore, zero-shot prediction results highlight its strong generalization capabilities. This framework offers a promising solution for bridging protein domain knowledge with natural language understanding, paving the way for transformative advancements in protein-related research.
1 Introduction
With the continuous advancement of biotechnology, gaining a deeper understanding of protein sequences, structures, and their functions has become increasingly important. To elucidate the complex mechanisms underlying protein-mediated physiological processes, researchers have focused on studying and predicting the physicochemical properties of proteins. The widespread availability of protein sequences and structures, enabled by novel high-throughput sequencing technologies and advanced structure prediction tools Jumper et al. (2021); Abramson et al. (2024), has greatly facilitated the exploration of intricate relationships between protein sequences, structures, and functions. In this context, protein Q&A systems have emerged as a research hotspot, aiming to help scientists and researchers efficiently access information about protein properties and functions Zhou et al. (2023). These systems are of considerable significance in fields such as drug discovery and disease research. Besides, just as some proteins mentioned in AF-Cluster Wayment-Steele et al. (2024), it indicates that even if the protein sequences are the same, they may present different 3D structures due to different environmental conditions, binding partners, or point mutations, thereby performing different biological functions.
Traditional approaches, such as those based on Gene Ontology (GO) classifications, predict protein functions by leveraging sequence information through convolutional models or sequence similarity-based computations, as exemplified by methods like DeepGOPlus Kulmanov and Hoehndorf (2020). Such models primarily utilize protein sequence information and are limited to classification tasks, which oversimplifies the complexity of protein functions. The recent rise of large language models (LLMs) has propelled protein Q&A systems into a cutting-edge research domain, allowing scientists to efficiently access and interpret critical information about protein properties and functions Zhou et al. (2023). These systems hold immense potential for transformative applications in areas such as drug discovery and disease research. To support this growing field, several datasets, including Mol-Instructions Fang et al. (2023) and UniProtQA Luo et al. (2023), have been developed, providing essential resources for advancing protein Q&A research.
An increasing number of models have fine-tuned LLMs using protein-specific Q&A datasets to generate contextually relevant text or answer questions related to protein structure and function. However, as highlighted in Table 1, existing approaches often face significant limitations, including a lack of multimodal information integration, limited flexibility, and high computational costs. These shortcomings restrict their ability to provide precise and contextually appropriate responses. For instance, FAPM Xiang et al. (2024) and InstructProtein Wang et al. (2023), despite incorporating generative Q&A functionalities, remain heavily influenced by classification-based frameworks. This often results in the generation of extraneous or irrelevant information, deviating from the intended scope of the answers and limiting their utility in addressing complex protein-related queries. BiomedGPT Luo et al. (2023) introduces a generative framework tailored for protein Q&A tasks but relies solely on sequence information, which significantly limits its predictive capabilities.
We compared the generation effects of large language models trained with protein structures and sequences, respectively, on different types of questions, as shown in Figure 1. The results show that when answering questions related to protein subcellular localization and domains or motifs, the generation effect of the model jointly trained with protein structures and sequences is significantly better than that trained only with protein sequences. Some following works such as ProtChatGPT Wang et al. (2024) and Evola Zhou et al. (2025) have tried to employ two separate modules for encoding sequence and structural information. However, such practice makes it difficult for protein sequence and structural features to interact, and significantly increases computational complexity. Moreover, no existing model has systematically examined the relative contributions of sequence and structural information to the performance of protein Q&A systems. This gap underscores the critical need for future research to evaluate how integrating both sequence and structural data can enhance the accuracy, robustness, and interpretability of protein-related Q&A.
As the sequence and structural features of proteins are one-to-one correspondence, we propose a novel framework called Prot2Chat that seamlessly incorporates the spatial structure and sequence information into LLMs. This approach seeks to enhance both the performance and applicability of protein Q&A systems. As illustrated in Figure 2, the overall architecture of our model comprises three primary components: the protein encoder, the protein-to-text adapter, and the large language model. Specifically, we extended the structure encoder ProteinMPNN Dauparas et al. (2022) to fuse sequence information during node initialization. The sequence features we introduced are inherent to the model, thus eliminating the need for additional training parameters. Then following the methodology of InstructPLM Qiu et al. (2024), a lightweight cross-attention layer was incorporated to align the protein encoder with a LoRA Hu et al. (2022) fine-tuned LLM. This approach preserves the primary sequence and structural information of proteins during the compression process. The processed protein structure and sequence data are provided as a soft prompt to the LLM. Leveraging the flexible text generation capabilities of the LLM, our model effectively follows residue-level protein instructions, significantly enhancing its utility in protein-related tasks.
To evaluate the model’s performance, we conducted comparative experiments on the Mol-Instructions and UniprotQA datasets. And we conducted zero-shot experiments using a part of the UniprotQA dataset to verify the generalization ability of the model. The generated answers were assessed using common metrics such as BLEU and ROUGE, as well as online KIMI Qin et al. (2024) scoring metrics and expert evaluations, ensuring the reliability and relevance of the outputs. Experimental results demonstrate that the model employing a unified encoder to integrate protein sequence and structural information outperforms those using sequence-only pretraining or ESM-based encoders Lin et al. (2022). Furthermore, the high consistency between expert evaluations and online KIMI assessments confirms the robustness of our model.
In summary, our contributions are as follows:
  1. 1.
    We extended the existing structure encoder ProteinMPNN to realize structure and sequence early fusion without the need for training.
  2. 2.
    Based on the protein encoder, we achieved a lightweight and effective protein large language model with only 93M training parameters.
  3. 3.
    We performed various systematic assessments across various evaluation datasets to validate our model’s generative and generalization capabilities. These included traditional metrics, large model evaluations online, and manual expert evaluations.
Figure 1: Comparison of Q&A results between Prot2Chat and LLaMA3-FT with sequence.
Figure 2: Model Structure of Prot2Chat. The red font represents the input, the snowflake represents freezing, and the fire represents the parameters to be trained.
2 Related work
Protein Representation Learning
Proteins are fundamental components of cells, essential for their biological activities and diverse functions. Previous studies on protein characterization have explored various methods to learn protein representations based on different forms of protein information. Protein sequences, often referred to as the ”language of life,” have been extensively studied using advanced natural language processing techniques. For instance, Tranception Notin et al. (2022) employs the Transformer Vaswani (2017) model to encode amino acid sequences, capturing relationships between residues and autoregressively reconstructing protein sequences from large-scale databases. Similarly, sequential modeling approaches such as ESM Rives et al. (2021); Lin et al. (2023); Hayes et al. (2024) leverage masked language modeling (MLM) to develop attention patterns that correspond to residue-residue contact maps, enhancing sequence-based protein representations. On the other hand, structure-based approaches Gligorijević et al. (2021); Achiam et al. (2023) directly indicates protein function and encodes geometric information of the protein for topology-sensitive tasks such as protein property prediction. Foldseek van Kempen et al. (2022) introduces the concept of a structural alphabet, encoding protein structures into a discrete representation space. Similarly, Saprot Su et al. (2023) introduces a structure-aware vocabulary, embedding structural information into model inputs to enhance representational capacity, achieving significant success in protein function prediction tasks.
Multi-Modal Alignment
Enabling large language models (LLMs) to understand additional modalities has been a rapidly evolving research direction, with notable examples including image-text models Tsimpoukelli et al. (2021); Li et al. (2023), video-text models like VideoLlama Zhang et al. (2023), audio-text models such as Macaw-LLM Lyu et al. (2023), and molecular-text models, including ReLM Shi et al. (2023), and MolTC Fang et al. (2024). This line of research was pioneered by advancements in visual language modeling (VLM), which has been successfully applied to tasks such as few-shot image classification, image captioning, and image Q&A Li et al. (2023); Alayrac et al. (2022). To enable LLMs to understand images, leading VLM methods adopt different strategies. Some, like BLIP-2 Li et al. (2023), employ nonlinear and expressive cross-modal projectors, while others, such as PaLM-E Driess et al. (2023), utilize visual encoders and fine-tune LLMs on multimodal datasets. These approaches have also been increasingly applied to protein analysis and understanding. For instance, Galactica Taylor et al. (2022) leverages scientific literature to model protein sequences and SMILES representations, enabling the model to interpret sequence properties. ProtNote Char et al. (2024), utilize two encoders to encode protein sequence information and text information, respectively, and Multilayer Perceptron (MLP) to fuse these inputs, achieving the goal of using text to achieve supervised and zero-shot protein functional prediction. ProteinChat Guo et al. (2023) further uses the protein structure and the corresponding description to model for the protein Q&A tasks. ProtChatGPT Wang et al. (2024) uses two independent modules to encode sequence and structural information respectively, and aligns them with natural language to explore protein functions. InstructBioMol Zhuang et al. (2024) integrates multimodal biomolecular inputs through a feature extraction module, allowing researchers to articulate design goals in natural language. InstructProtein Wang et al. (2023) employs a supervised training framework based on instruction generation using knowledge graphs, enabling bidirectional generation between natural language and protein language. This model can predict protein functional descriptions and generate protein sequences based on natural language prompts. Prot2Text Abdine et al. (2024) combines graph neural networks (GNNs) with LLMs to predict protein functions in free-text form. Similarly, ProtT3 Liu et al. (2024) provides an efficient protein-to-text generation framework by integrating protein language models (PLMs) with LLMs. FAPM Xiang et al. (2024) utilizes a contrastive learning framework to implement a generative-like Q&A model, while BiomedGPT Luo et al. (2023) bridges the gap between biological and human natural language through a multimodal generative pre-trained transformer (GPT). This allows users to ”communicate” with biological modalities using free-text queries, facilitating interactions with biological data in natural language.
Model Input Output
Prot2Text Sequence Structured Text
ProtNote Sequence&Description Function&Probability
ProtChatGPT Structure&Sequence&Question Free Text
Evola Structure&Sequence&Question Free Text
InstructProtein Sequence&Question Structured Text
FAPM Sequence Structured Text
BioMedGPT Sequence&Question Free Text
Prot2Chat [Structure+Sequence]&Question Free Text
Table 1: Comparison of input-output of protein Q&A models, ”Description” is a text related to protein functions, such as the description of Gene Ontology (GO) terms, ”[Structure + Sequence]” represents the early fusion of structure and sequence.
Dataset Train Valid Test
Mol-Instructions 404640 16859 11072
UniProtQA 25820 1075 6734
Table 2: The count and division of the datasets we used.
Prompt
Please select the sentence that is closest and most accurate in meaning to the Target sentence and rank sentences A, B, and C accordingly, with the first place meaning the most identical and accurate in meaning to the Target. Only provide the answer, for example, ‘B A C’ or ‘A C B’, etc.
Target: {Ground Truth} A: {Model1 Generated Text} B: {Model2 Generated Text} C: {Model3 Generated Text}
Table 3: Prompt to KIMI for evaluation
Model Bleu-2 Rouge-1 Rouge-2 Rouge-L
LLaMA3-8B-Instruct 4.84 23.07 4.31 15.36
LLaMA3-FT with sequence 6.42 24.50 6.32 17.03
BioMedGPT-LM-10B 1.02 10.93 1.57 7.84
KIMI (zero-shot) 4.79 22.21 4.62 14.70
KIMI (few-shot) 12.05 31.21 11.38 24.18
Prot2Chat 33.25 54.88 35.26 47.90
w/o fine-tuned LLM 12.87 30.89 14.67 26.81
w/o Protein sequence 31.88 52.90 33.43 46.17
Table 4: Result in Mol-instructions protein oriented dataset, the best performances are marked in bold.
Model Bleu-2 Rouge-1 Rouge-2 Rouge-L
LLaMA3-8B-Instruct 2.41 13.66 1.10 9.07
BioMedGPT-LM-10B 5.80 15.88 7.32 14.68
Prot2Chat (zero-shot) 2.11 13.78 1.51 9.64
Prot2Chat (fine-tuned) 5.87 15.65 9.00 14.99
Table 5: Result in UniprotQA dataset, the best performances are marked in bold.
Model 1st 2nd 3rd Avg.
Prot2Chat 591 386 22 1.43
BioMedGPT-LM-10B 218 214 567 2.35
LLaMA3-FT with sequence 190 399 410 2.22
Table 6: Results of KIMI evaluation, ”Avg.” represents the average ranking.
Model 1st 2nd 3rd Avg.
Prot2Chat 454 115 78 1.42
BioMedGPT-LM-10B 157 203 287 2.20
LLaMA3-FT with sequence 36 329 282 2.38
Table 7: Results of expert manual evaluation, ”Avg.” represents the average ranking.
Inputs LLaMA3-FT with sequence Prot2Chat Ground Truth
[Uncaptioned image] [O58787]Using the provided structure and amino acid sequence of a protein, estimate its primary location within the cell: Based on the provided amino acid sequence, it appears that the protein is a membrane protein, likely localized to the endoplasmic reticulum, Golgi apparatus, and/or plasma membrane. The presence of multiple transmembrane domains and the distribution of hydrophobic and hydrophilic residues suggest that this protein is an integral membrane protein. Additionally, the sequence contains several potential phosphorylation sites and potential N-glycosylation sites, which may be involved in protein-protein interactions and/or post-translational modification. The protein characterized by the structure and amino acid sequence demonstrates rRNA binding and is implicated in the translation. Its subcellular localization is primarily within the cytosol. Upon analyzing the given protein structure and sequence, it appears that the protein is likely to be localized in the cytosol compartment of the cell.
[Uncaptioned image] [O58736]Using the protein structure and the following protein sequence as input, assess its potential function by comparing it to known proteins with similar sequences or structures: The provided protein sequence is approximately 240 amino acids long and shows no significant similarity to known proteins in the database. However, it can be analyzed for potential functional regions and motifs using various bioinformatic tools. Here are some observations and predictions: 1. **Transmembrane regions**: The sequence contains several regions that are predicted to be transmembrane, including the following: * 24-45 amino acids (ILLLVGMPGSGKGEVAKAFRRRGIPVI) * 63-82 amino acids (MGGDAIREEAEKRGIPKTPEGLKYVSLKVREE) With the provided structure and amino acid sequence, the protein is expected to display adenylate kinase activity and participate in the phosphorylation, ATP biosynthetic process, phosphorylation. The protein with the structure and amino acid sequence has been analyzed, and its structure suggests a potential involvement in ATP binding.
Table 8: Case study. The bold font represents the key points of the answers, the blue font represents the correct answers of the key points, and the red font represents the incorrect answers of the key points.
3 Method
3.1 Overview of Prot2Chat
As illustrated in Figure 2, our model leverages the 3D coordinate structure information of protein residue atoms along with the protein sequence. The protein encoder processes this information to generate a protein embedding representation, which is then transformed into a text-aligned protein prompt via the protein-text adapter. This prompt, combined with the input question, serves as the input to the large language model (LLM), ultimately producing the output answer.
3.2 Sequence and Structure Fused Protein Encoder
[…]
$h_{E}$ $=\text{Linear}(E)$ (1)
$h_{V}$ $=\text{Zeros}(n,Dc)$ (2)
$h_{V},h_{E}$ $=\text{Encoderlayers}(h_{V},h_{E})$ (3)
$h_{V}$ $=\text{Decoderlayers}(h_{V},h_{E})$ (4)
where Linear, Encoderlayers and Decoderlayers are all modules in ProteinMPNN, $E$ is the structual features and $n$ is the number of protein residues and $Dc$ is the embedding dimension of the protein encoder.
We find that a variant version of ProteinMPNN additionally contains the Embedding module for protein sequences $S$, and the Decoderlayer is sequence-aware, which changes Equation 4 to:
$h_{V}=\text{Decoderlayers}(h_{V},h_{E},\text{Embedding}(S))$ (5)
To be specific, the Decoderlayer combines the node embeddings, the output features of the encoder, and the sequence embeddings to construct the context features for decoding through neighborhood feature aggregation. Using the message passing mechanism, the hidden state $h_{V}$ is continuously updated through multi-layer stacking to simulate the dynamic generation process of the decoder.
On this basis, we further initialized $h_{V}$ with the off-the-shelf sequence embedding to make the sequence and structural information early fused. Specifically, Equation 2 is replaced by:
$h_{V}=\text{Embedding}(S)$ (6)
In particular, all of the model weights we used are from ProteinMPNN, with no need for additional training. The final protein node vector ${h}_{V}$ is used as the protein feature for the protein-text adapter. Inspired by InstructPLM Qiu et al. (2024), we concatenate the protein representations encoded by nine released ProteinMPNN models. As a result, $Dc=128*9=1152$.
3.3 Protein-Text Adapter
We implemented a protein-text adapter to semantically align the information obtained by the protein encoder with natural language, and provided both types of information as input to the large model simultaneously. The feature $h_{V}$ obtained from the modified ProteinMPNN model is output through the linear projection layer, positional encoding and the cross-attention mechanism. Specifically, the input protein feature first passes through a linear projection layer $W_{\text{proj}}$ to transform the input features to the target output dimension $D_{\text{o}}$:
$X_{\text{proj}}=W_{\text{proj}}\cdot h_{V}+PE$ (7)
where $PE$ is the Dynamic Positional Encoding Vaswani (2017) used to capture the positional information of amino acids.
The complete protein encoding is too long. We adopt the idea of BLIP-2 to extract the important semantic features from it. In particular, we introduce $n_{\text{q}}$ learnable queries $\boldsymbol{Q}\in\mathbb{R}^{n_{\text{q}}\times D_{\text{o}}}$ and also apply $PE$ to them. Then the multi-head attention layer is adapted to capture key protein information based on queries:
$A^{k}=\text{softmax}\left(\frac{Q^{k}K^{k\top}}{\sqrt{D_{k}}}\right)V^{k}$ (8)
where $Q^{k}=\boldsymbol{Q}W_{Q}^{k}$, $K^{k}=X_{\text{proj}}W_{K}^{k}$ and $V^{k}=X_{\text{proj}}W_{V}^{k}$ stand for queries, keys and values. Among them, $W_{Q}^{k},W_{K}^{k},W_{V}^{k}$ are related parameters, $k$ is the head index. The multi-head attention outputs are then concatenated across heads and linearly transformed to the final protein prompt:
$\boldsymbol{X_{\text{protein}}}=\text{Concat}(\boldsymbol{A}^{1},\boldsymbol{A}^{2},\cdots,\boldsymbol{A}^{M})\boldsymbol{W}_{out}$ (9)
3.4 LLM Decoder
We combine $X_{\text{protein}}$ representing protein information with the text question and input it into the existing LLM to obtain the response.
$\text{response}=\text{LLM}(X_{\text{protein}},\text{question})$ (10)
To improve domain adaptability, we fine-tuned the LLM with LoRA while training the adapter. The number of adapter training parameters is 89,702,400, while the number of LLaMA3 training parameters is 3,407,872. The total number of training parameters of the Prot2Chat is 93M. Meanwhile, the number of training parameters of BiomedGPT is 3B, that of ProtT3 is approximately 2B, the range of trainable parameters of Prot2Text is from 256M to 898M, and that of FAPM is 188M. This also shows the advantage of our model in significantly reducing the computational cost. Besides, we adpot the CrossEntropyLoss function as common.
4 Experiment
4.1 Datasets
We evaluated the performance of our model using the Mol-Instructions Fang et al. (2023) and UniProtQA Luo et al. (2023) datasets. Mol-Instructions is a comprehensive instruction dataset specifically designed for the biomolecular domain, comprising both human-constructed and machine-generated data. For our experiments, we utilized the protein-oriented instruction subset, which is primarily derived from entries in the UniProtKB/Swiss-Prot database UniProt Consortium (2018). This subset encompasses tasks such as predicting protein domains, functions, and activities. Our model was predominantly trained on this dataset.
Additionally, we employed a portion of the UniProtQA dataset, introduced by BiomedGPT Luo et al. (2023), to assess the generalization capability of our model through zero-shot evaluation. Fine-tuning was also performed on this dataset to further validate performance. UniProtQA consists of textual descriptions of proteins, including their functions and properties. It was curated by UniProt UniProt Consortium (2018) and includes four types of questions related to protein function, formal name, protein family, and subcellular location. The detailed sizes of our dataset splits are provided in Table 2.
To incorporate structural and sequence information, we retrieved the corresponding PDB files for proteins listed in Swiss-Prot. These PDB files, combined with the associated questions, serve as the input to Prot2Chat.
4.2 Baselines
We introduced the following baselines that represent protein information using sequence.
  • LLaMA3: LLaMA3 Dubey et al. (2024) series has attracted a lot of research attention due to its outstanding capabilities in the general domain. We use LLaMA3-8B-Instruct and input the protein sequence and question to achieve zero-shot.
  • LLaMA3-FT: We also fine-tuned LLaMA3-8B-Instruct using textual protein sequence information and used this model as a reference model.
  • BioMedGPT-10B: BioMedGPT is a domain-specific LLM trained on a large selection of corpora of human scientific knowledge. The model encodes the protein sequence with the ESM-3B Lin et al. (2022) and uses BioMedGPT-LM-7B as the decoder to generate the response.
  • KIMI: We used the online large model KIMI as a control.
In order to make a fair comparison, we used the same input template for each model through prompt engineering, so that the model could output the answer as we instructed. Some other open-source models in related fields, such as FAPM, Prot2Text, and InstructProtein, were not included in the comparison because these models complete protein Q&A tasks based on classification, or there are differences in the formats of the model’s inputs and outputs, as shown in Table 1. Therefore, it is hard to conduct comparisons.
4.3 Model Setting
We jointly trained the protein adapter and LLaMA3 on the Mol-Instructions dataset, employing full training for the adapter and LoRA (Low-Rank Adaptation) fine-tuning for LLaMA3. Inspired by BLIP-2 Li et al. (2023), we configured the adapter with 256 queries. For the LoRA setup, we set the rank $r$ to 8, LoRA alpha to 16, and targeted the ”q_proj” and ”v_proj” modules, with a LoRA dropout rate of 0.1. The adapter comprises 89,702,400 trainable parameters, while LLaMA3 has 3,407,872 trainable parameters.
We utilized the Adam optimizer and implemented gradient accumulation to optimize the training process. The initial learning rate was set to $10^{-4}$, with a batch size of 2 and a maximum context length of 1024 tokens, which includes both the question and answer text. The model was fine-tuned for 2 epochs, with each training session requiring approximately 1600 hours on an NVIDIA RTX 3090 GPU.
4.4 Evaluation Metrics
To evaluate the effectiveness of the model’s text generation, we employed performance metrics including BLEU Papineni et al. (2002) and ROUGE Lin (2004). Additionally, we utilized the prompt templates shown in Table 3 to simultaneously input the target text and the text generated by different models into the online KIMI model. This allowed us to determine which generated answer was closer to the target text, thereby assisting in the evaluation of the model’s output quality. The results, as shown in Table 6, demonstrate that our model outperforms other comparative models overall. Furthermore, we conducted expert manual evaluations, where professional biology PhDs ranked the responses of different models based on their alignment with the target text. The results of these evaluations (Table 7) are consistent with those obtained using KIMI.
4.5 Main Results
The results of the assessment are shown in Table 4 and Table 5. The results show that there is a significant modal gap between the protein sequence and the natural language, and it is incomprehensible to directly input the protein sequence as text and problem into the large language model, resulting in the disordered and meaningless response of the model. Aligning protein language with human language is an effective solution to solve this problem, and we have observed that the performance of the model trained with the addition of structural information is significantly better than that of other models. This suggests that we can directly use the structure and sequence information of proteins to help language models better understand and generate technical terms through feature space alignment. Establishing a link between proteins and natural language can provide guidance for researchers in their research on unknown proteins. The results of our model’s prediction on the UniprotQA Luo et al. (2023) dataset verify the generalization performance of our model. Further illustrating the role of our model in protein Q&A tasks.
4.6 Ablation Experiments
As shown in Table 4, we conducted a comparative analysis to evaluate the impact of incorporating protein structure during training, thereby highlighting the importance of structural information for protein understanding. We also examined the effect of fine-tuning LLMs, revealing that training adapters alone was insufficient for achieving optimal performance. Furthermore, by comparing the results obtained from training with sequence text alone versus training with both sequence and structural information, we observed that LLMs struggle to effectively interpret protein sequence information presented in textual form. These findings underscore the necessity of integrating protein structural information alongside a unified multimodal alignment module, which enables LLMs to better understand and generate domain-specific technical terms.
4.7 Case Study
As shown in Table 8, the results we obtained using the protein Q&A test dataset indicate that the model trained with the addition of protein structure information generates responses significantly better than the model trained with protein sequence. Further prove that it is difficult for LLM to discover the corresponding protein functions and effects only through the text information of protein sequences, and thus draw the conclusion that the protein structure is vital for the comprehensive understanding of proteins.
5 Conclusion
This paper introduces a novel protein Q&A model that addresses the limitations of traditional protein Q&A tasks by leveraging the modified ProteinMPNN and an adapter module to integrate protein structure and sequence information with natural language text. By harnessing the flexible generation capabilities of large language models (LLMs), our approach effectively bridges the gap between protein data and textual understanding. Experimental results demonstrate that the model trained with structural information significantly outperforms other baseline models, underscoring the critical role of protein structure in protein understanding and analysis. Furthermore, the results validate the effectiveness and strong generalization ability of our model in protein Q&A tasks.

### v2

@@
-Prot2Chat: Protein LLM with Early Fusion of Sequence and Structure
+
 Abstract
-Proteins play a pivotal role in living organisms, yet understanding their functions presents significant challenges, including the limited flexibility of classification-based methods, the inability to effectively leverage spatial structural information, and the lack of systematic evaluation metrics for protein Q&A systems.
-To address these limitations, we propose Prot2Chat, a novel framework that integrates multimodal protein representations with natural language through a unified module, enabling large language model (LLM)-driven answer generation.
-Our model incorporates a modified ProteinMPNN encoder, which encodes protein sequence and structural information in a unified manner, a protein-text adapter with cross-attention mechanisms, and a LLaMA3 decoder.
-To optimize training efficiency, we freeze the encoder and employ LoRA techniques for the decoder.
-We conducted experiments on two datasets, both automated metrics and expert evaluations demonstrate the superior performance of our model.
-Furthermore, zero-shot prediction results highlight its strong generalization capabilities.
-This framework offers a promising solution for bridging protein domain knowledge with natural language understanding, paving the way for transformative advancements in protein-related research.
+Motivation: Proteins are of great significance in living organisms.
+However, understanding their functions encounters numerous challenges, such as insufficient integration of multimodal information, a large number of training parameters, limited flexibility of classification-based methods, and the lack of systematic evaluation metrics for protein Q&A systems.
+To tackle these issues, we propose the Prot2Chat framework.
+Results: We modified ProteinMPNN to encode protein sequence and structural information in a unified way.
+We used a large language model (LLM) to encode questions into vectors and developed a protein-text adapter to compress protein information into virtual tokens based on these vectors, achieving the early fusion of text and protein information.
+Finally, the same LLM reads the virtual tokens and the questions to generate answers.
+To optimize training efficiency, we froze the encoder and employed Low-Rank Adaptation (LoRA) techniques for the LLM.
+Experiments on two datasets show that both automated metrics and expert evaluations demonstrate the superior performance of our model, and zero-shot prediction results highlight its generalization ability.
+The models and codes are available at https://github.com/wangzc1233/Prot2Chat.
+Contact: zqcao@suda.edu.cn or wangzc025@163.com
+keywords:
+Protein Q&A, Early-Fusion, LLM
 1 Introduction
-With the continuous advancement of biotechnology, gaining a deeper understanding of protein sequences, structures, and their functions has become increasingly important.
-To elucidate the complex mechanisms underlying protein-mediated physiological processes, researchers have focused on studying and predicting the physicochemical properties of proteins.
-The widespread availability of protein sequences and structures, enabled by novel high-throughput sequencing technologies and advanced structure prediction tools Jumper et al.
-(2021); Abramson et al.
-(2024), has greatly facilitated the exploration of intricate relationships between protein sequences, structures, and functions.
-In this context, protein Q&A systems have emerged as a research hotspot, aiming to help scientists and researchers efficiently access information about protein properties and functions Zhou et al.
-(2023).
-These systems are of considerable significance in fields such as drug discovery and disease research.
-Besides, just as some proteins mentioned in AF-Cluster Wayment-Steele et al.
-(2024), it indicates that even if the protein sequences are the same, they may present different 3D structures due to different environmental conditions, binding partners, or point mutations, thereby performing different biological functions.
-Traditional approaches, such as those based on Gene Ontology (GO) classifications, predict protein functions by leveraging sequence information through convolutional models or sequence similarity-based computations, as exemplified by methods like DeepGOPlus Kulmanov and Hoehndorf (2020).
-Such models primarily utilize protein sequence information and are limited to classification tasks, which oversimplifies the complexity of protein functions.
-The recent rise of large language models (LLMs) has propelled protein Q&A systems into a cutting-edge research domain, allowing scientists to efficiently access and interpret critical information about protein properties and functions Zhou et al.
-(2023).
-These systems hold immense potential for transformative applications in areas such as drug discovery and disease research.
-To support this growing field, several datasets, including Mol-Instructions Fang et al.
-(2023) and UniProtQA Luo et al.
-(2023), have been developed, providing essential resources for advancing protein Q&A research.
-An increasing number of models have fine-tuned LLMs using protein-specific Q&A datasets to generate contextually relevant text or answer questions related to protein structure and function.
-However, as highlighted in Table 1, existing approaches often face significant limitations, including a lack of multimodal information integration, limited flexibility, and high computational costs.
-These shortcomings restrict their ability to provide precise and contextually appropriate responses.
-For instance, FAPM Xiang et al.
-(2024) and InstructProtein Wang et al.
-(2023), despite incorporating generative Q&A functionalities, remain heavily influenced by classification-based frameworks.
-This often results in the generation of extraneous or irrelevant information, deviating from the intended scope of the answers and limiting their utility in addressing complex protein-related queries.
-BiomedGPT Luo et al.
-(2023) introduces a generative framework tailored for protein Q&A tasks but relies solely on sequence information, which significantly limits its predictive capabilities.
-We compared the generation effects of large language models trained with protein structures and sequences, respectively, on different types of questions, as shown in Figure 1.
-The results show that when answering questions related to protein subcellular localization and domains or motifs, the generation effect of the model jointly trained with protein structures and sequences is significantly better than that trained only with protein sequences.
-Some following works such as ProtChatGPT Wang et al.
-(2024) and Evola Zhou et al.
-(2025) have tried to employ two separate modules for encoding sequence and structural information.
-However, such practice makes it difficult for protein sequence and structural features to interact, and significantly increases computational complexity.
+With the continued advancement of biotechnology, gaining a deeper understanding of protein sequences, structures, and their functions has become increasingly critical.
+Protein function is fundamentally governed by its amino acid sequence, which encodes the structural and biochemical features necessary for activity.
+However, even identical protein sequences can adopt distinct three-dimensional conformations under varying environmental conditions or subtle mutations, thereby exhibiting diverse biological functions, as exemplified by cases reported in AFcluster wayment2024predicting .
+The widespread availability of protein sequence and structure data—enabled by high-throughput sequencing technologies and powerful structure prediction tools abramson2024accurate has significantly advanced our understanding of the physicochemical properties of proteins.
+In parallel, significant research efforts have focused on investigating and predicting protein functions, where deep learning models have employed sequence or structure representations as input to predict function-related properties such as Gene Ontology (GO) terms kulmanov2020deepgoplus ; su2023saprot .
+The recent emergence of large language models (LLMs) has propelled protein question answering (Q&A) systems at the forefront of computational biology research zhou2023survey .
+The related systems enable researchers to efficiently access and interpret key information regarding protein properties and functions.
+This holds great potential for transformative applications in fields such as drug discovery and disease research smith2025funcfetch .
+To support this growing field, several datasets, such as Mol-Instructions fang2023mol and UniProtQA luo2023biomedgpt , have been introduced, providing essential resources for advancing protein Q&A systems.
+An increasing number of models have fine-tuned LLMs on protein-specific Q&A datasets to generate contextually relevant text or to answer questions related to protein structure and function zhou2025decoding .
+However, as summarized in Table 1, existing methods face several critical limitations, including inadequate integration of multimodal information, restricted flexibility, high training parameter volume, and a lack of systematic evaluation metrics.
+These challenges constrain the models’ ability to generate accurate and context-aware responses.
+For instance, although models such as FAPM xiang2024fapm and InstructProtein wang2023instructprotein incorporate generative Q&A capabilities, they remain heavily dependent on classification-based paradigms.
+This often leads to the production of unexpected content, deviating from the intended scope of inquiry and reducing the models’ effectiveness in handling complex protein-related questions.
+BiomedGPT luo2023biomedgpt enables users to interact with biological data through natural language.
+Nonetheless, it relies exclusively on sequence-level information, which substantially limits its predictive capacity.
+ProtChatGPT wang2024protchatgpt and Evola zhou2025decoding attempt to address this limitation by employing separate modules for encoding sequence and structural information.
+However, this modular design hinders effective interaction between sequence and structural features and significantly increases computational complexity.
 Moreover, no existing model has systematically examined the relative contributions of sequence and structural information to the performance of protein Q&A systems.
-This gap underscores the critical need for future research to evaluate how integrating both sequence and structural data can enhance the accuracy, robustness, and interpretability of protein-related Q&A.
-As the sequence and structural features of proteins are one-to-one correspondence, we propose a novel framework called Prot2Chat that seamlessly incorporates the spatial structure and sequence information into LLMs.
+Figure 1: Prot2Chat can assist human in understanding protein information and achieve cross-modal information communication.
+Among them, ‘$<$Soft Prompt$>$‘ is the prompt obtained by our model through fusing protein structure, sequence, and text information, which helps the LLM generate more valuable answers.
+As the sequence and structural features of proteins are in inseparable correspondence, both jointly determine the properties of proteins.
+We propose a novel framework named Prot2Chat.
+This framework can seamlessly integrate the spatial structure and sequence information of proteins into LLMs.
+Meanwhile, by incorporating text information during the processing of protein information, the LLM is enabled to understand the question, so as to achieve question-based compression of protein information, and ultimately realize the early fusion of text, structure, and sequence.
 This approach seeks to enhance both the performance and applicability of protein Q&A systems.
-As illustrated in Figure 2, the overall architecture of our model comprises three primary components: the protein encoder, the protein-to-text adapter, and the large language model.
-Specifically, we extended the structure encoder ProteinMPNN Dauparas et al.
-(2022) to fuse sequence information during node initialization.
-The sequence features we introduced are inherent to the model, thus eliminating the need for additional training parameters.
-Then following the methodology of InstructPLM Qiu et al.
-(2024), a lightweight cross-attention layer was incorporated to align the protein encoder with a LoRA Hu et al.
-(2022) fine-tuned LLM.
-This approach preserves the primary sequence and structural information of proteins during the compression process.
-The processed protein structure and sequence data are provided as a soft prompt to the LLM.
+To be specific, we introduce protein sequence information during the initialization of nodes features in ProteinMPNN dauparas2022robust and early-fuse the protein sequence information in the protein feature representation.
+Then inspired by BLIP-2 li2023blip , an adapter module was incorporated to align the protein encoder with a LoRA hu2022lora fine-tuned LLM.
+In addition, we use the LLM to encode the question text and add it to the adapter, integrating it into the above-mentioned compression process to further achieve the early-fusion of protein and text information, aiming to improve the accuracy and robustness of protein Q&A.
+The processed protein and text data are provided as a soft prompt to the LLM.
 Leveraging the flexible text generation capabilities of the LLM, our model effectively follows residue-level protein instructions, significantly enhancing its utility in protein-related tasks.
-To evaluate the model’s performance, we conducted comparative experiments on the Mol-Instructions and UniprotQA datasets.
-And we conducted zero-shot experiments using a part of the UniprotQA dataset to verify the generalization ability of the model.
-The generated answers were assessed using common metrics such as BLEU and ROUGE, as well as online KIMI Qin et al.
-(2024) scoring metrics and expert evaluations, ensuring the reliability and relevance of the outputs.
-Experimental results demonstrate that the model employing a unified encoder to integrate protein sequence and structural information outperforms those using sequence-only pretraining or ESM-based encoders Lin et al.
-(2022).
-Furthermore, the high consistency between expert evaluations and online KIMI assessments confirms the robustness of our model.
+As shown in Figure 1, the answering effect of our model can provide effective insights for human to explore proteins in a Q&A manner familiar to them.
+To evaluate the model’s performance, we conducted comparative experiments on the Mol-Instructions fang2023mol and UniProtQA luo2023biomedgpt datasets, and performed zero-shot experiments on a subset of UniProtQA to validate generalization.
+We used common metrics like BLEU and ROUGE, as well as online KIMI qin2024mooncake scoring metrics and expert evaluations, ensuring the reliability and relevance of the answers.
+Experimental results show that Prot2Chat which integrates protein sequences and structures using a unified encoder and fuses text information at an early stage outperforms those using sequence-only pretraining, ESM-based encoders lin2022language , and multi-encoders.
+Besides, its effectiveness far outperforms that of late-fusion, which means that text information is only utilized when the LLM generates answers.
+These findings highlight the effectiveness of early-fusion strategies for integrating multimodal protein representations.
+Furthermore, the high consistency between expert evaluations and online KIMI assessments confirms the robustness of Prot2Chat.
 In summary, our contributions are as follows:
+  1. [1.]
   2. 1.
-    We extended the existing structure encoder ProteinMPNN to realize structure and sequence early fusion without the need for training.
+    We extended the existing structure encoder ProteinMPNN to realize structure and sequence early-fusion without the need for training.
   3. 2.
-    Based on the protein encoder, we achieved a lightweight and effective protein large language model with only 93M training parameters.
+    Based on this protein encoder, we designed a text-aware protein-text adapter module.
+By re-utilizing the LLM, we achieved question-based compression of protein information, early-fusion of text and protein information, and implemented a lightweight and efficient protein LLM with only 109 million training parameters.
   4. 3.
@@
 These included traditional metrics, large model evaluations online, and manual expert evaluations.
-Figure 1: Comparison of Q&A results between Prot2Chat and LLaMA3-FT with sequence.
-Figure 2: Model Structure of Prot2Chat. The red font represents the input, the snowflake represents freezing, and the fire represents the parameters to be trained.
 2 Related work
-Protein Representation Learning
+2.1 Protein Representation Learning
 Proteins are fundamental components of cells, essential for their biological activities and diverse functions.
@@
 Protein sequences, often referred to as the ”language of life,” have been extensively studied using advanced natural language processing techniques.
-For instance, Tranception Notin et al.
-(2022) employs the Transformer Vaswani (2017) model to encode amino acid sequences, capturing relationships between residues and autoregressively reconstructing protein sequences from large-scale databases.
-Similarly, sequential modeling approaches such as ESM Rives et al.
-(2021); Lin et al.
-(2023); Hayes et al.
-(2024) leverage masked language modeling (MLM) to develop attention patterns that correspond to residue-residue contact maps, enhancing sequence-based protein representations.
-On the other hand, structure-based approaches Gligorijević et al.
-(2021); Achiam et al.
-(2023) directly indicates protein function and encodes geometric information of the protein for topology-sensitive tasks such as protein property prediction.
-Foldseek van Kempen et al.
-(2022) introduces the concept of a structural alphabet, encoding protein structures into a discrete representation space.
-Similarly, Saprot Su et al.
-(2023) introduces a structure-aware vocabulary, embedding structural information into model inputs to enhance representational capacity, achieving significant success in protein function prediction tasks.
-Multi-Modal Alignment
-Enabling large language models (LLMs) to understand additional modalities has been a rapidly evolving research direction, with notable examples including image-text models Tsimpoukelli et al.
-(2021); Li et al.
-(2023), video-text models like VideoLlama Zhang et al.
-(2023), audio-text models such as Macaw-LLM Lyu et al.
-(2023), and molecular-text models, including ReLM Shi et al.
-(2023), and MolTC Fang et al.
-(2024).
-This line of research was pioneered by advancements in visual language modeling (VLM), which has been successfully applied to tasks such as few-shot image classification, image captioning, and image Q&A Li et al.
-(2023); Alayrac et al.
-(2022).
-To enable LLMs to understand images, leading VLM methods adopt different strategies.
-Some, like BLIP-2 Li et al.
-(2023), employ nonlinear and expressive cross-modal projectors, while others, such as PaLM-E Driess et al.
-(2023), utilize visual encoders and fine-tune LLMs on multimodal datasets.
-These approaches have also been increasingly applied to protein analysis and understanding.
-For instance, Galactica Taylor et al.
-(2022) leverages scientific literature to model protein sequences and SMILES representations, enabling the model to interpret sequence properties.
-ProtNote Char et al.
-(2024), utilize two encoders to encode protein sequence information and text information, respectively, and Multilayer Perceptron (MLP) to fuse these inputs, achieving the goal of using text to achieve supervised and zero-shot protein functional prediction.
-ProteinChat Guo et al.
-(2023) further uses the protein structure and the corresponding description to model for the protein Q&A tasks.
-ProtChatGPT Wang et al.
-(2024) uses two independent modules to encode sequence and structural information respectively, and aligns them with natural language to explore protein functions.
-InstructBioMol Zhuang et al.
-(2024) integrates multimodal biomolecular inputs through a feature extraction module, allowing researchers to articulate design goals in natural language.
-InstructProtein Wang et al.
-(2023) employs a supervised training framework based on instruction generation using knowledge graphs, enabling bidirectional generation between natural language and protein language.
+For instance, Tranception notin2022tranception employs the Transformer vaswani2017attention model to encode amino acid sequences, capturing relationships between residues and autoregressively reconstructing protein sequences from large-scale databases.
+Similarly, sequential modeling approaches such as ESM lin2023evolutionary leverage masked language modeling (MLM) to develop attention patterns that correspond to residue-residue contact maps, enhancing sequence-based protein representations.
+On the other hand, structure-based approaches gligorijevic2021structure directly indicate protein function and encode geometric information of the protein for topology-sensitive tasks such as protein property prediction.
+Foldseek van2022foldseek introduces the concept of a structural alphabet, encoding protein structures into a discrete representation space.
+Similarly, Saprot su2023saprot introduces a structure-aware vocabulary, embedding structural information into model input to enhance representational capacity, achieving significant success in protein function prediction tasks.
+2.2 Multimodal Alignment
+Enabling LLMs to understand extra modalities has been a fast-evolving research area, with notable examples including image-text models li2023blip , video-text models like VideoLlama zhang2023video , audio-text models such as Macaw-LLM lyu2023macaw , and molecular-text models like MolTC fang2024moltc .
+This line of research was pioneered by advancements in visual language modeling (VLM), to enable LLMs to understand images, leading VLM methods adopt different strategies.
+Some, like BLIP-2 li2023blip , employ nonlinear and expressive cross-modal projectors, while others, like PaLM-E driess2023palm , utilize visual encoders and fine-tune LLMs on multimodal datasets.
+These approaches have also been increasingly applied to protein analysis.
+For instance, Galactica taylor2022galactica leverages scientific literature to model protein sequences and SMILES representations, enabling the model to interpret sequence properties.
+Some of these methods attempt to adopt a Q&A approach that is closer to human habits to explore and understand the functions of proteins.
+ProtNote char2024protnote , utilize two encoders to encode protein sequence information and text information, respectively, and Multilayer Perceptron (MLP) to fuse these inputs, achieving the goal of using text to achieve supervised and zero-shot protein functional prediction.
+ProtChatGPT wang2024protchatgpt uses two independent modules to encode sequence and structural information respectively, and aligns them with natural language to explore protein functions.
+InstructProtein wang2023instructprotein employs a supervised training framework based on instruction generation using knowledge graphs, enabling bidirectional generation between natural language and protein language.
 This model can predict protein functional descriptions and generate protein sequences based on natural language prompts.
-Prot2Text Abdine et al.
-(2024) combines graph neural networks (GNNs) with LLMs to predict protein functions in free-text form.
-Similarly, ProtT3 Liu et al.
-(2024) provides an efficient protein-to-text generation framework by integrating protein language models (PLMs) with LLMs.
-FAPM Xiang et al.
-(2024) utilizes a contrastive learning framework to implement a generative-like Q&A model, while BiomedGPT Luo et al.
-(2023) bridges the gap between biological and human natural language through a multimodal generative pre-trained transformer (GPT).
+Prot2Text abdine2024prot2text combines graph neural networks (GNNs) with LLMs to predict protein functions in free-text form.
+FAPM xiang2024fapm utilizes a contrastive learning framework to implement a generative-like Q&A model, while BiomedGPT luo2023biomedgpt bridges the gap between biological and human natural language through a multimodal generative pre-trained transformer (GPT).
 This allows users to ”communicate” with biological modalities using free-text queries, facilitating interactions with biological data in natural language.
+Figure 2: Model Structure of Prot2Chat.
+The red font represents the input, the snowflake represents freezing, and the flame represents the parameters to be trained.
+We obtain the embedding with multi-dimensional feature fusion of proteins from protein structure and sequence information through an encoder.
+Meanwhile, we get the question vector from the question text.
+Then, we conduct early-fusion and alignment of this vector with protein information to obtain the soft prompt.
+Finally, we input the soft prompt along with the question text into the LLM to get the answer.
+Table 1: Comparison of input-output of protein Q&A models, ”Description” is a text related to protein functions, such as the description of Gene Ontology (GO) terms, ”[Structure+Sequence+Question]” represents the early-fusion of structure, sequence and text.
 Model Input Output
@@
 BioMedGPT Sequence&Question Free Text
-Prot2Chat [Structure+Sequence]&Question Free Text
-Table 1: Comparison of input-output of protein Q&A models, ”Description” is a text related to protein functions, such as the description of Gene Ontology (GO) terms, ”[Structure + Sequence]” represents the early fusion of structure and sequence.
+Prot2Chat [Structure+Sequence+Question]&Question Free Text
+Table 2: The count and division of the datasets we used.
 Dataset Train Valid Test
@@
 UniProtQA 25820 1075 6734
-Table 2: The count and division of the datasets we used.
-Prompt
-Please select the sentence that is closest and most accurate in meaning to the Target sentence and rank sentences A, B, and C accordingly, with the first place meaning the most identical and accurate in meaning to the Target.
-Only provide the answer, for example, ‘B A C’ or ‘A C B’, etc.
-Target: {Ground Truth} A: {Model1 Generated Text} B: {Model2 Generated Text} C: {Model3 Generated Text}
-Table 3: Prompt to KIMI for evaluation
+Table 3: Result in Mol-instructions protein oriented dataset, the best performances are marked in bold.
 Model BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L
@@
 BioMedGPT-LM-10B 1.02 10.93 1.57 7.84
+Evola-10B 8.69 29.09 8.41 20.04
 KIMI (Zero-shot) 4.79 22.21 4.62 14.70
 KIMI (Few-shot) 12.05 31.21 11.38 24.18
-Prot2Chat 33.25 54.88 35.26 47.90
+Prot2Chat 35.85 57.21 38.09 50.51
+Table 4: Results of the ablation experiments in Mol-instructions dataset. The best performance is marked in bold.
+Model BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L
+Prot2Chat 35.85 57.21 38.09 50.51
 w/o Fine-tuned LLM 12.87 30.89 14.67 26.81
 w/o Protein sequence 31.88 52.90 33.43 46.17
-Table 4: Result in Mol-instructions protein oriented dataset, the best performances are marked in bold.
+w/o Early-fusing text 33.25 54.88 35.26 47.90
+Table 5: Result in UniProtQA dataset, the best performances are marked in bold.
 Model BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L
@@
 BioMedGPT-LM-10B 5.80 15.88 7.32 14.68
-Prot2Chat (zero-shot) 2.11 13.78 1.51 9.64
-Prot2Chat (fine-tuned) 5.87 15.65 9.00 14.99
-Table 5: Result in UniprotQA dataset, the best performances are marked in bold.
-Model 1st 2nd 3rd Avg.
-Prot2Chat 591 386 22 1.43
-BioMedGPT-LM-10B 218 214 567 2.35
-LLaMA3-FT with sequence 190 399 410 2.22
+Prot2Chat (Zero-shot) 2.23 14.73 1.73 10.10
+Prot2Chat (Fine-tuned) 6.72 15.71 9.25 15.57
 Table 6: Results of KIMI evaluation, ”Avg.” represents the average ranking.
-Model 1st 2nd 3rd Avg.
-Prot2Chat 454 115 78 1.42
-BioMedGPT-LM-10B 157 203 287 2.20
-LLaMA3-FT with sequence 36 329 282 2.38
+Model 1st 2nd 3rd 4th Avg.
+Prot2Chat 386 242 13 9 1.45
+BioMedGPT-LM-10B 26 57 326 241 3.20
+LLaMA3-FT with sequence 18 22 252 358 3.46
+Evola-10B 220 329 59 42 1.88
 Table 7: Results of expert manual evaluation, ”Avg.” represents the average ranking.
+Model 1st 2nd 3rd 4th Avg.
+Prot2Chat 359 266 17 8 1.49
+BioMedGPT-LM-10B 28 58 334 230 3.18
+LLaMA3-FT with sequence 11 21 239 379 3.52
+Evola-10B 252 305 60 33 1.81
+Table 8: Case study.
+The bold font represents the key points of the answers, the blue font represents the correct answers of the key points, and the red font represents the incorrect answers of the key points.
 Inputs LLaMA3-FT with sequence Prot2Chat Ground Truth
-[Uncaptioned image] [O58787]Using the provided structure and amino acid sequence of a protein, estimate its primary location within the cell: Based on the provided amino acid sequence, it appears that the protein is a membrane protein, likely localized to the endoplasmic reticulum, Golgi apparatus, and/or plasma membrane.
-The presence of multiple transmembrane domains and the distribution of hydrophobic and hydrophilic residues suggest that this protein is an integral membrane protein.
-Additionally, the sequence contains several potential phosphorylation sites and potential N-glycosylation sites, which may be involved in protein-protein interactions and/or post-translational modification.
-The protein characterized by the structure and amino acid sequence demonstrates rRNA binding and is implicated in the translation.
-Its subcellular localization is primarily within the cytosol.
-Upon analyzing the given protein structure and sequence, it appears that the protein is likely to be localized in the cytosol compartment of the cell.
 [Uncaptioned image] [O58736]Using the protein structure and the following protein sequence as input, assess its potential function by comparing it to known proteins with similar sequences or structures: The provided protein sequence is approximately 240 amino acids long and shows no significant similarity to known proteins in the database.
@@
 Here are some observations and predictions: 1.
-**Transmembrane regions**: The sequence contains several regions that are predicted to be transmembrane, including the following: * 24-45 amino acids (ILLLVGMPGSGKGEVAKAFRRRGIPVI) * 63-82 amino acids (MGGDAIREEAEKRGIPKTPEGLKYVSLKVREE) With the provided structure and amino acid sequence, the protein is expected to display adenylate kinase activity and participate in the phosphorylation, ATP biosynthetic process, phosphorylation.
+Transmembrane regions: The sequence contains several regions that are predicted to be transmembrane, including the following: 24-45 amino acids (ILLLVGMPGSGKGEVAKAFRRRGIPVI) 63-82 amino acids (MGGDAIREEAEKRGIP KTPEGLKYVSLKVREE) With the provided structure and amino acid sequence, the protein is expected to display adenylate kinase activity and participate in the phosphorylation, ATP biosynthetic process, phosphorylation.
 The protein with the structure and amino acid sequence has been analyzed, and its structure suggests a potential involvement in ATP binding.
-Table 8: Case study.
-The bold font represents the key points of the answers, the blue font represents the correct answers of the key points, and the red font represents the incorrect answers of the key points.
 3 Method
 3.1 Overview of Prot2Chat
-As illustrated in Figure 2, our model leverages the 3D coordinate structure information of protein residue atoms along with the protein sequence.
-The protein encoder processes this information to generate a protein embedding representation, which is then transformed into a text-aligned protein prompt via the protein-text adapter.
-This prompt, combined with the input question, serves as the input to the large language model (LLM), ultimately producing the output answer.
+As illustrated in Figure 2, the overall architecture of our model comprises three components: the protein encoder, the text-aware protein-text adapter, and the LLM.
+Our model leverages the 3D coordinate structure information of protein residue atoms along with the protein sequence and natural language question text.
+The protein encoder processes protein information to generate a protein embedding representation, which is then transformed into a text-aligned protein prompt via the text-aware protein-text adapter, with question text information being fused simultaneously.
+This prompt, combined with the input question, serves as the input to the large language model, ultimately producing the output answer.
 3.2 Sequence and Structure Fused Protein Encoder
-[…]
+We modified the structure encoder ProteinMPNN to realize the sequence and structure early-fusion.
+ProteinMPNN is originally dedicated to design protein sequences based on backbone structures.
+The input is structural information $E$, namely the 3D coordinates of protein residue atoms ($N$: amino nitrogen, $C\alpha$: alpha carbon, $C$: carbonyl carbon and $O$: carbonyl oxygen).
+Then protein features $h_{V}$ are obtained through the encoder layer and the decoder layer.
 $h_{E}$ $=\text{Linear}(E)$ (1)
@@
 To be specific, the Decoderlayer combines the node embeddings, the output features of the encoder, and the sequence embeddings to construct the context features for decoding through neighborhood feature aggregation.
-Using the message passing mechanism, the hidden state $h_{V}$ is continuously updated through multi-layer stacking to simulate the dynamic generation process of the decoder.
-On this basis, we further initialized $h_{V}$ with the off-the-shelf sequence embedding to make the sequence and structural information early fused. Specifically, Equation 2 is replaced by:
+Using the message passing mechanism, the hidden state $h_{V}$ is continuously updated through multilayer stacking to simulate the dynamic generation process of the decoder.
+On this basis, we further initialized $h_{V}$ with the off-the-shelf sequence embedding to make the sequence and structural information early-fused. Specifically, Equation 2 is replaced by:
 $h_{V}=\text{Embedding}(S)$ (6)
 In particular, all of the model weights we used are from ProteinMPNN, with no need for additional training.
-The final protein node vector ${h}_{V}$ is used as the protein feature for the protein-text adapter.
-Inspired by InstructPLM Qiu et al.
-(2024), we concatenate the protein representations encoded by nine released ProteinMPNN models.
+The final protein node vector ${h}_{V}$ is used as the protein feature embedding for the text-aware protein-text adapter.
+Inspired by InstructPLM qiu2024instructplm , we concatenate the protein representations encoded by nine released ProteinMPNN models.
 As a result, $Dc=128*9=1152$.
-3.3 Protein-Text Adapter
-We implemented a protein-text adapter to semantically align the information obtained by the protein encoder with natural language, and provided both types of information as input to the large model simultaneously.
-The feature $h_{V}$ obtained from the modified ProteinMPNN model is output through the linear projection layer, positional encoding and the cross-attention mechanism.
-Specifically, the input protein feature first passes through a linear projection layer $W_{\text{proj}}$ to transform the input features to the target output dimension $D_{\text{o}}$:
+3.3 Text-Aware Protein-Text Adapter
+We implemented a text-aware protein-text adapter to semantically align the information obtained from the protein encoder with natural language, incorporating question text information simultaneously.
+Subsequently, this fused information is provided as a soft prompt to the LLM.
+The protein embedding $h_{V}$ obtained from the modified ProteinMPNN model is output through the linear projection layer, positional encoding and the cross-attention mechanism.
+During the above-mentioned process, text information is fused simultaneously, as shown in Figure  3.
+Figure 3: Detail of the Text-Aware Protein-Text Adapter.
+The adapter takes protein embedding and question vector as inputs.
+It integrates text information at an early stage through a set of learnable queries.
+Subsequently, using these queries, the protein embedding is used as keys/values for cross-attention calculations.
+This cross-attention module compresses the protein information into a fixed length and captures the key protein information according to the queries infused with question text information, resulting in soft prompt to assist the LLM in generating more accurate responses.
+To be specific, the input protein feature first passes through a linear projection layer $W_{\text{proj}}$, the aim is to transform the input features to the target output dimension $D_{\text{o}}$:
 $X_{\text{proj}}=W_{\text{proj}}\cdot h_{V}+PE$ (7)
-where $PE$ is the Dynamic Positional Encoding Vaswani (2017) used to capture the positional information of amino acids.
+where $PE$ is the Dynamic Positional Encoding vaswani2017attention used to capture the positional information of amino acids.
+Meanwhile, we utilize the question text and the LLM to obtain the hidden state $Qh_{t}$ of the last text token as the question vector, so as to achieve prompt compression and carry out the early-fusion of text and protein information, enabling the model to perceive question text information earlier:
+$Qh_{\text{t}}=\text{LLM}(\text{question}).hiddenstate$ (8)
 The complete protein encoding is too long.
 We adopt the idea of BLIP-2 to extract the important semantic features from it.
-In particular, we introduce $n_{\text{q}}$ learnable queries $\boldsymbol{Q}\in\mathbb{R}^{n_{\text{q}}\times D_{\text{o}}}$ and also apply $PE$ to them.
-Then the multi-head attention layer is adapted to capture key protein information based on queries:
+In particular, we introduce $n_{\text{q}}$ learnable queries $\boldsymbol{Q}\in\mathbb{R}^{n_{\text{q}}\times D_{\text{o}}}$, incorporate the question vector $Qh_{t}$ obtained from the question text, and also apply the positional encoding (PE) to them.
+Then the multi-head cross-attention layer is adapted to capture key protein information based on queries:
 $A^{k}=\text{softmax}\left(\frac{Q^{k}K^{k\top}}{\sqrt{D_{k}}}\right)V^{k}$ (9)
@@
 Among them, $W_{Q}^{k},W_{K}^{k},W_{V}^{k}$ are related parameters, $k$ is the head index.
-The multi-head attention outputs are then concatenated across heads and linearly transformed to the final protein prompt:
+Then, the outputs of multi-head cross-attention are concatenated across heads and linearly transformed by ${W}_{out}$ into the final protein prompt integrated with text information, which serves as the soft prompt for the LLM, where $M$ is the number of attention heads:
 $\boldsymbol{X_{\text{protein}}}=\text{Concat}(\boldsymbol{A}^{1},\boldsymbol{A}^{2},\cdots,\boldsymbol{A}^{M})\boldsymbol{W}_{out}$ (10)
 3.4 LLM Decoder
-We combine $X_{\text{protein}}$ representing protein information with the text question and input it into the existing LLM to obtain the response.
+We combine $X_{\text{protein}}$, which integrates text and protein information, with the text question, and then input this combination into the existing LLM to obtain a response.
 $\text{response}=\text{LLM}(X_{\text{protein}},\text{question})$ (11)
 To improve domain adaptability, we fine-tuned the LLM with LoRA while training the adapter.
-The number of adapter training parameters is 89,702,400, while the number of LLaMA3 training parameters is 3,407,872.
-The total number of training parameters of the Prot2Chat is 93M.
-Meanwhile, the number of training parameters of BiomedGPT is 3B, that of ProtT3 is approximately 2B, the range of trainable parameters of Prot2Text is from 256M to 898M, and that of FAPM is 188M.
+The number of adapter training parameters is 106,483,712, while the number of LLaMA3 training parameters is 3,407,872.
+The total number of training parameters of the Prot2Chat is 109M.
+Meanwhile, the number of training parameters of BiomedGPT-10B is 3B, that of Evola-10B is approximately 1.7B, the range of trainable parameters of Prot2Text is from 256M to 898M, and that of FAPM is 188M.
 This also shows the advantage of our model in significantly reducing the computational cost.
@@
 4.1 Datasets
-We evaluated the performance of our model using the Mol-Instructions Fang et al.
-(2023) and UniProtQA Luo et al.
-(2023) datasets.
-Mol-Instructions is a comprehensive instruction dataset specifically designed for the biomolecular domain, comprising both human-constructed and machine-generated data.
-For our experiments, we utilized the protein-oriented instruction subset, which is primarily derived from entries in the UniProtKB/Swiss-Prot database UniProt Consortium (2018).
+We evaluated the performance of our model using the Mol-Instructions fang2023mol and UniProtQA luo2023biomedgpt datasets.
+For our experiments, we utilized the protein-oriented instruction subset in Mol-Instructions, which is primarily derived from entries in the UniProtKB/Swiss-Prot database uniprot2018uniprot .
 This subset encompasses tasks such as predicting protein domains, functions, and activities.
 Our model was predominantly trained on this dataset.
-Additionally, we employed a portion of the UniProtQA dataset, introduced by BiomedGPT Luo et al.
-(2023), to assess the generalization capability of our model through zero-shot evaluation.
+Additionally, we employed a portion of the UniProtQA dataset, introduced by BiomedGPT luo2023biomedgpt , to assess the generalization capability of our model through zero-shot evaluation.
 Fine-tuning was also performed on this dataset to further validate performance.
 UniProtQA consists of textual descriptions of proteins, including their functions and properties.
-It was curated by UniProt UniProt Consortium (2018) and includes four types of questions related to protein function, formal name, protein family, and subcellular location.
 The detailed sizes of our dataset splits are provided in Table 2.
+It is worth noting that there are significant differences in the text style and length between the UniProtQA and Mol-Instructions datasets.
+This directly leads to large differences in the traditional metrics of the subsequent evaluation test results.
 To incorporate structural and sequence information, we retrieved the corresponding PDB files for proteins listed in Swiss-Prot.
@@
 We introduced the following baselines that represent protein information using sequence.
-  • LLaMA3: LLaMA3 Dubey et al.
-(2024) series has attracted a lot of research attention due to its outstanding capabilities in the general domain.
+  • LLaMA3: LLaMA3 dubey2024llama series has attracted a lot of research attention due to its outstanding capabilities in the general domain.
 We use LLaMA3-8B-Instruct and input the protein sequence and question to achieve zero-shot.
   • LLaMA3-FT: We also fine-tuned LLaMA3-8B-Instruct using textual protein sequence information and used this model as a reference model.
-  • BioMedGPT-10B: BioMedGPT is a domain-specific LLM trained on a large selection of corpora of human scientific knowledge.
-The model encodes the protein sequence with the ESM-3B Lin et al.
-(2022) and uses BioMedGPT-LM-7B as the decoder to generate the response.
+  • BioMedGPT-LM-10B: BioMedGPT luo2023biomedgpt is a domain-specific LLM trained on a large selection of corpora of human scientific knowledge.
+The model encodes the protein sequence with the ESM-3B lin2022language and uses BioMedGPT-LM-7B as the decoder to generate the response.
+  • Evola-10B: Evola zhou2025decoding is trained on hundreds of millions of protein Q&A pairs and an AI-generated dataset of 150 billion words.
+It adopts Retrieval-Augmented Generation (RAG) to integrate external knowledge and performs excellently in generating accurate, detailed, and context-relevant answers about protein functions, providing expert-level insights for proteomics and functional genomics research.
+Here, we use its model with 10B parameters as the comparative model.
   • KIMI: We used the online large model KIMI as a control.
@@
 Therefore, it is hard to conduct comparisons.
+In addition, the ProtChatGPT model was not included in the comparison because accessible resources for this model could not be found.
 4.3 Model Setting
-We jointly trained the protein adapter and LLaMA3 on the Mol-Instructions dataset, employing full training for the adapter and LoRA (Low-Rank Adaptation) fine-tuning for LLaMA3.
-Inspired by BLIP-2 Li et al.
-(2023), we configured the adapter with 256 queries.
+We jointly trained the text-aware protein-text adapter and LLaMA3 on the Mol-Instructions dataset, employing full training for the adapter and LoRA fine-tuning for LLaMA3.
+Inspired by InstructPLM qiu2024instructplm , we configured the adapter with 256 queries and integrate text information on each query.
 For the LoRA setup, we set the rank $r$ to 8, LoRA alpha to 16, and targeted the ”q_proj” and ”v_proj” modules, with a LoRA dropout rate of 0.1.
-The adapter comprises 89,702,400 trainable parameters, while LLaMA3 has 3,407,872 trainable parameters.
+The adapter comprises 106,483,712 trainable parameters, while LLaMA3 has 3,407,872 trainable parameters.
 We utilized the Adam optimizer and implemented gradient accumulation to optimize the training process.
@@
 4.4 Evaluation Metrics
-To evaluate the effectiveness of the model’s text generation, we employed performance metrics including BLEU Papineni et al.
-(2002) and ROUGE Lin (2004).
-Additionally, we utilized the prompt templates shown in Table 3 to simultaneously input the target text and the text generated by different models into the online KIMI model.
-This allowed us to determine which generated answer was closer to the target text, thereby assisting in the evaluation of the model’s output quality.
-The results, as shown in Table 6, demonstrate that our model outperforms other comparative models overall.
+To evaluate the effectiveness of the model’s text generation, we employed performance metrics including BLEU papineni-etal-2002-bleu and ROUGE lin-2004-rouge .
+Additionally, the target text and the texts generated by different models are simultaneously input into the online KIMI model, using the following prompts: Please select the sentence that is closest and most accurate in meaning to the Target sentence and rank sentences A, B, C and D accordingly, with the first place meaning the most identical and accurate in meaning to the Target.
+Only provide the answer, for example, ‘B D A C’ or ‘A C D B’, etc.
+Target: {Ground Truth} A: {Model1 Generated Text} B: {Model2 Generated Text} C: {Model3 Generated Text} D: {Model4 Generated Text} This allowed us to determine which generated answer was closer to the target text, thereby assisting in the evaluation of the model’s output quality.
 Furthermore, we conducted expert manual evaluations, where professional biology PhDs ranked the responses of different models based on their alignment with the target text.
-The results of these evaluations (Table 7) are consistent with those obtained using KIMI.
 4.5 Main Results
-The results of the assessment are shown in Table 4 and Table 5.
-The results show that there is a significant modal gap between the protein sequence and the natural language, and it is incomprehensible to directly input the protein sequence as text and problem into the large language model, resulting in the disordered and meaningless response of the model.
-Aligning protein language with human language is an effective solution to solve this problem, and we have observed that the performance of the model trained with the addition of structural information is significantly better than that of other models.
+The results of the assessment are shown in Table 3 to Table 7.
+The BLEU-2 and ROUGE-L scores of Prot2Chat exceed those of models such as BiomedGPT and Evola.
+The results show that there is a significant modal gap between the protein sequence and the natural language, and it is incomprehensible to directly input the protein sequence as text and problem into the LLM, resulting in the disordered and meaningless response of the model.
+Aligning protein language with human language is an effective way to solve this problem.
+Moreover, we have observed that the performance of the model trained through the early-fusion of structural information and sequences is significantly superior to that of other models leveraging either sequences or structures.
 This suggests that we can directly use the structure and sequence information of proteins to help language models better understand and generate technical terms through feature space alignment.
-Establishing a link between proteins and natural language can provide guidance for researchers in their research on unknown proteins.
-The results of our model’s prediction on the UniprotQA Luo et al.
-(2023) dataset verify the generalization performance of our model.
+In addition, the early-fusion of text information and protein information also demonstrates a significant effect.
+Unification of multimodal protein information and establishment of a connection between proteins and natural language can provide guidance for researchers in their study of unknown proteins.
+The results of our model’s prediction on the UniProtQA luo2023biomedgpt dataset verify the generalization performance of our model.
+On the other hand, the evaluation results of online large language models, as shown in Table 6, demonstrate that our model outperforms other comparative models overall and the results of expert manual evaluations (Table 7) are consistent with those obtained using KIMI.
 Further illustrating the role of our model in protein Q&A tasks.
 4.6 Ablation Experiments
-As shown in Table 4, we conducted a comparative analysis to evaluate the impact of incorporating protein structure during training, thereby highlighting the importance of structural information for protein understanding.
-We also examined the effect of fine-tuning LLMs, revealing that training adapters alone was insufficient for achieving optimal performance.
-Furthermore, by comparing the results obtained from training with sequence text alone versus training with both sequence and structural information, we observed that LLMs struggle to effectively interpret protein sequence information presented in textual form.
-These findings underscore the necessity of integrating protein structural information alongside a unified multimodal alignment module, which enables LLMs to better understand and generate domain-specific technical terms.
+As shown in Table 4, we carried out a comparative analysis to evaluate the impact of incorporating protein structure information during the training process.
+Comparing with the results of the model trained with only the sequence further highlights the importance of structural information for understanding proteins.
+We also studied the effect of fine-tuning LLMs.
+The results show that training only the adapter is insufficient to achieve optimal performance.
+By comparing the results obtained from training in the cases of early-fusion and non-early-fusion of text and protein information, we found that the early-fusion of information plays a non-negligible role in the understanding of proteins and the generation of large language models, and its effect is significantly better than that of late-fusion.
+These findings emphasize the necessity of integrating protein structure information and a unified multimodal alignment module, which enables large language models to better understand and generate domain-specific technical terms.
 4.7 Case Study
-As shown in Table 8, the results we obtained using the protein Q&A test dataset indicate that the model trained with the addition of protein structure information generates responses significantly better than the model trained with protein sequence.
-Further prove that it is difficult for LLM to discover the corresponding protein functions and effects only through the text information of protein sequences, and thus draw the conclusion that the protein structure is vital for the comprehensive understanding of proteins.
+As shown in Table 8, the results obtained from our use of the protein Q&A test dataset indicate that the responses generated by the model trained by adding protein structure and early-fusing text information are significantly superior to those of other trained models (taking LLaMA3-FT as an example here).
+This further demonstrates that protein structure is of crucial importance for the comprehensive understanding of proteins, the early-fusion and alignment of multimodal information plays a non-negligible role in the understanding of proteins and the generation of large language models.
 5 Conclusion
-This paper introduces a novel protein Q&A model that addresses the limitations of traditional protein Q&A tasks by leveraging the modified ProteinMPNN and an adapter module to integrate protein structure and sequence information with natural language text.
+This paper introduces a novel protein Q&A model that addresses the limitations of traditional protein Q&A tasks by leveraging the modified ProteinMPNN and a text-aware adapter module to integrate protein structure and sequence information with natural language text.
 By harnessing the flexible generation capabilities of large language models (LLMs), our approach effectively bridges the gap between protein data and textual understanding.
-Experimental results demonstrate that the model trained with structural information significantly outperforms other baseline models, underscoring the critical role of protein structure in protein understanding and analysis.
-Furthermore, the results validate the effectiveness and strong generalization ability of our model in protein Q&A tasks.
+Experimental results demonstrate that the model trained by adding protein structure and early-fusing text information significantly outperforms other baseline models.
+This not only highlights the crucial role of protein structure in protein understanding and analysis but also proves the effectiveness of the early-fusion of protein information and text.
+Furthermore, the results validate the effectiveness and generalization ability of our model in protein Q&A tasks.